# Product-oriented Product–Process–Resource Asset Network and its Representation in AutomationML for Asset Administration Shell

### v1

@@ Abstract
Current products, especially in the automotive sector, pose complex technical systems having a multi-disciplinary mechatronic nature. Industrial standards supporting system engineering and production typically (i) address the production phase only, but do not cover the complete product life cycle, and (ii) focus on production processes and resources rather than the products themselves. The presented approach is motivated by incorporating impacts of end-of-life phase of the product life cycle into the engineering phase. This paper proposes a modelling approach coming up from the product-process-resource (PPR) modeling paradigm. It combines requirements on (i) respecting the product structure as a basis for the model, and (ii) it incorporates repairing, remanufacturing, or upcycling within cyber-physical production systems. The proposed model called PoPAN should accompany the product during the entire life cycle as a digital shadow encapsulated within the Asset Administration Shell of a product. To facilitate adoption of the proposed paradigm, the paper also proposes serialization of the model in the AutomationML data format. The model is demonstrated on a use-case for disassembling electric vehicle batteries to support their remanufacturing for stationary battery applications.
@@ I Introduction
The significant environmental impact of manufacturing industries, underscores the importance of addressing sustainability in the context of recycling and remanufacturing of products [1]. Although the term ”sustainability” is increasingly used in marketing, leading to skepticism and negative perceptions. Distinguishing genuine sustainability efforts from mere marketing strategies is essential. In this paper, we introduce a formalization to model products and production processes, which has a potential to genuinely contribute to sustainability specifically at the end of a product lifecycle.
@@
PoPAN offers a structured description of a product, incorporating its components (represented as products), processes, and resources. This approach enables the design of recycling and remanufacturing processes for products that require disassembly for efficient recycling. Furthermore, PoPAN can function as a digital shadow [4], accompanying the product throughout its entire lifecycle and encapsulating all relevant information. It serves as a record of the product’s evolution, including any modifications or alterations it undergoes during its journey, such as missing components like screws. An additional advantage of employing PoPAN for recycling lies in the data collection aspect. Sustainability initiatives begin at the product design phase, where prioritizing recyclability is paramount. By gathering relevant data from recycling processes, manufacturers can analyze and utilize this information to enhance the recyclability of future product designs. This iterative process leads to more sustainable designs and facilitates easier recycling, contributing to overall environmental sustainability efforts another advantage of a digital twin/PoPAN for recycling is the collection of data from the process.
To facilitate the adoption of this paradigm, the paper also proposes the serialization of the model in the AutomationML [5] data format. This standardized format enhances interoperability and ease of integration within existing industrial frameworks and systems. An important benefit of AutomationML is the advancing integration with the Asset Administration Shell (AAS) [6]. AAS is a set of standards and recommendations to provide interoperability in industrial systems. AAS is typically used on the resource level (e.g., a motor or a robot). On the contrary, the AAS is used for the product level in the presented approach utilizing PoPAN. Moreover, AAS is used for the entire product life-cycle in this approach, including not only the manufacturing phase, but also operations, maintenance and repairing, as well as decommissioning/recycling. The paper [7] highlights the use of AAS in the context of the product life cycle management.
@@
  • RQ1: How can we adapt the PAN model to primary reflect the product structure rather than the production process with resource structure?
@@
  • RQ3: To improve adoption of the proposed approach by industry and academia, how can be the proposed model represented in the AutomationML data format?
The proposed modeling approach is demonstrated on a case study on disassembling electric vehicle (EV) batteries for recycling and remanufacturing [8].
@@ II State of the Art
The proposed approach is built on the top of 3 main domains including (i) Asset Administration Shell, (ii) AutomationML data format, and (iii) Model-Driven Engineering and Product-Process-Resource Asset Network. These research areas are described in details in the following subsections.
@@ II-A Asset Administration Shell
One of the key standards in the Industry 4.0 initiative is the Asset Administration Shell (AAS) [9, 10]. AAS was introduced by the Platform Industry 4.0 as a promising approach to enable interoperability in various industries. It facilitates data sharing between value chain partners, standardizes data security, and establishes technology-neutral semantic standards, making it crucial to achieving Industry 4.0 goals. The AAS enables diverse communication channels and applications, serving as a bridge between physical objects and the digital world [6].
The AAS is a digital representation of an asset and comprises multiple submodels that represent a specific aspect or set of characteristics of an asset. Within a submodel, various submodel elements are defined. Submodel elements are the individual components or entities that make up the submodel. They represent specific properties, operations, parameters, references, files, or other relevant information associated with the asset. The purpose of submodels is to provide a structured approach to organizing and managing the elements within an AAS. By grouping related elements, it becomes easier to define and enforce standards for specific sets of information. By combining the AAS with the reference architecture model Industry 4.0 (RAMI 4.0) [11], an asset becomes an I4.0 Component.
@@ II-B AutomationML
Engineering phase of multi-disciplinary mechatronic systems requires data exchange across software tools belonging to diverse engineering domains (including electric planning, mechanic design, maintenance, etc.). Engineering data can be efficiently captured and exchanged in the data format AutomationML, which is a standardized extension of XML for interdisciplinary exchange of planning data for production systems and processes. The data format AutomationML is open, platform-independent standard known as IEC 62714. Detailed information about this standard including various best practices and application recommendations, how to model specific information can be found on the Web pages of the AutomationML association [^1] Online: https://www.automationml.org/.
@@
  • CAEX (IEC 62424) – Topology a hierarchy of objects, properties and relationships among objects
@@
The basis of the object model in AutomationML is CAEX, which is intended for capturing engineering information in object-oriented form. Each object can have attributes and references to other objects (i.e., internal links or mapping objects, and references via object identifiers). It can also include links to other information, represented in COLLADA or PLCopen XML. In CAEX, model is represented via the following four basic corner-stones. Role Class Libraries are intended to define object semantics. Each object can have more than one roles. Interface Class Libraries provide specifications of object interfaces. Classes of objects are modeled as objects of System Unit Class Libraries. Each system unit class poses a generic object prototype, which can be instantiated when describing a specific system. For better interoperability, semantic mapping to terminology defined as ECLASS based on IEC 61360 can be used in a standardized way. The system description itself is represented in AutomationML/CAEX by instance hierarchy. In the tree structure behind the instance hierarchy root element, Internal Elements and their sub-elements are defined. Internal elements in the most cases reference their prototypes in the frame of System Unit Class, supported roles from Role Class Libraries and interfaces from the interface class libraries. Topological relationships are represented as Internal Links.
AutomationML is one of the most complex data formats for system engineering, which is suitable for a wide range of tasks, including the representation of the Asset Administration Shell [12, 13]. Such an AutomationML representation can be imported to the AASX format [14] in the tool AASX Package Explorer [^2] Online: https://github.com/eclipse-aaspe.
@@ II-C Model-Driven Engineering and Product-Process-Resource Asset Network
Model-driven engineering (MDE) highlights the role of models as core artifacts for engineering and integration development. Similarly like object-orientation in object-oriented programming, where everything is represented by objects (with specific attributes and methods), in model-driven engineering, everything is a model [15]. Initial efforts in MDE started with the use of Meta Object Facility and the Unified Modeling Language (UML) [16]. MDE is highly efficient for translating information between various representations (i.e., domain models or data formats) [17, 18]; and it provides a suitable support for multi-disciplinary engineering projects [19]. One of the plausible models to support the model-driven engineering is a product-process-resource model [20], categorizing artifacts relevant for industrial production into these three categories including mappings or links among these categories.
The Product-process-resource asset network (PAN) [3] is a graph-based formalism for expressing relationships among products, production processes, and resources. In the domain of Cyber-Physical Production Systems (CPPSs), dependencies among products, processes, and resources are often left implicit. The PAN addresses this gap by providing a structured framework for explicitly articulating and overseeing the dependencies among product, process, and resource assets. The PAN modeling is suitable for capturing and sharing knowledge among engineers and respective engineering tools as well as for identification of changes during engineering processes of multi-disciplinary system engineering projects. Usage of PAN optimizes efficiency of engineering processes and it enhances operations by modeling manufacturing processes, tracking resource usage, and monitoring asset performance. Through data integration, visualization, and analysis, PAN enables informed decision-making [21].
@@ III Product-oriented Product–Process–Resource Asset Network (PoPAN)
Figure 1: The proposed PoPAN model for a generic product (on the right), which has been created on the basis of the Product Structure model (on the left). The PoPAN graph is substantiated by products, which have assigned processes, and the edges (their types, orientation, and arrow types) have the crucial role.
@@
Recycling and remanufacturing products pose challenges due to limited engineering data availability. While manufacturing processes are well-specified and documented, the absence of detailed data for product in its end-of-life phase complicates recycling and remanufacturing efforts. Leveraging the PAN approach, which has proven efficiency in product manufacturing, presents a promising solution. However, traditional PAN methods describe one specific assembly process, which may not capture the entire variability of assembly and disassembly operations. To address this limitation, this paper proposes extending the PAN paradigm to structurally describe products. By doing so, multiple disassembly processes can be generated from a single description, accommodating the diverse disassembly options often encountered in practice. This structural representation enhances flexibility and efficiency in creating recycling and remanufacturing processes.
@@
To create a product description that offers a comprehensive understanding of the product’s structure, including the specific processes necessary for both assembly and disassembly, along with the resources required for each process, the initial step is to construct a Product Structure graph depicting dependencies between product’s components (see Fig. 1 on the left). Since the PPR approach is used, to preserve the name of elements, the product’s components are depicted also as products. Within this framework, three types of products are distinguished (1) Elementary Product: a product that cannot be further disassembled, (2) SubProduct: a product that is considered elementary in terms of assembly/disassembly procedure, but it is composed of several parts and can be further disassembled, and (3) Fastener Product: product that serves as a fastener for other products (e.g., screws). The graph delineates two types of edges connecting individual products, forming the overall structure. The solid line edge illustrates physical continuity between products and the dashed line edge on the other hand forms a structural continuity between the products, i. e., dependencies that indicate certain products cannot be assembled or disassembled until others are. Moreover, both types of edges end with arrows on both ends representing assembly/disassembly direction. Open arrows represent assembly direction, full arrows represent disassembly direction and Initial Product and Last Product are highlighted for better orientation in the graph. For example, when disassembling Product 3, it is essential to consider all three structural dependencies. Considering the orientation by the full arrows, Product 1 and Product 2 are not blocking Product 3 from being disassembled. However, the Last Product is blocking the disassembly of Product 3, meaning that the Last Product must be disassembled before Product 3. Additionally, frequently used terms final product and semi-product in the sense of a classic PAN are highlighted to show intermediate steps in an assemble operation. However, the PoPAN approach diverges from relying on these terms.
@@
Expanding the Product Structure graph with the PAN/PPR approach involves incorporating processes and resources alongside products. This structured visualization facilitates understanding the relationships between products and corresponding processes, which are supported by the required resources. As depicted in Fig. 1 on the right, in the extended Product Structure graph called Product-oriented Product–Process–Resource Asset Network (PoPAN), each product is connected with a corresponding process by a simple line edge or by a simple line edge with question mark above it. The question mark indicates conditional linking between the product and process, where the process sequence might not correspond with structure of a product, which is then resolved by red dotted line edge determining the sequence of the processes. Lastly, the graph contains green circle marks with a letter which depict edges connecting processes with their required resources. As with the previous Product Structure graph, the arrows form the sense of orientation in the graph and the connections between products, processes and resources form the edges.
@@
    Upon arriving at Fastener Product i, the conditional linking product to process edge is examined.
@@
    Returning to Fastener Product i, the physical continuity of products edge is verified in the correct direction, leading to Product 1.
@@
    Upon reaching Product 1, the structural continuity of products edge is examined, indicating the Product 3 is to be assembled subsequent to Product 1.
@@
    Fastener Process ii is noted with Fastener Product i simultaneously with Resource 1.
As outlined from the preceding steps, when assembling a product the initial edge to examine, if present, is the structural continuity of products edge. This edge determines whether the product can be assembled before other products. Once the feasibility of assembly/disassembly is established, the next edges to consider are the linking products to process edge and conditional linking products to process edge, which assign processes to products. Additionally, for the conditional linking products to process edge, the sequence of process edge is presented which ensures the processes are sequenced correctly. Afterward, the linking product to process edge is followed to assign resource to a product after which the physical continuity of products edge can be taken to continue to the next product where the whole process is repeated.
These rules, which specify the order in which edges should be traversed along with the assembly/disassembly direction, enable PoPAN to determine the correct path for creating assembly/disassembly operations. This capability facilitates efficient design of recycling or remanufacturing processes of products.

### v2

@@ Abstract
Current products, especially in the automotive sector, pose complex technical systems having a multi-disciplinary mechatronic nature. Industrial standards supporting system engineering and production typically (i) address the production phase only, but do not cover the complete product life cycle, and (ii) focus on production processes and resources rather than the products themselves. The presented approach is motivated by incorporating the impacts of the end-of-life phase of the product life cycle into the engineering phase. This paper proposes a modeling approach coming up from the Product-Process-Resource (PPR) modeling paradigm. It combines requirements on (i) respecting the product structure as a basis for the model, and (ii) incorporates repairing, remanufacturing, or upcycling within cyber-physical production systems. The proposed model called PoPAN should accompany the product during the entire life cycle as a digital shadow encapsulated within the Asset Administration Shell of a product. To facilitate the adoption of the proposed paradigm, the paper also proposes serialization of the model in the AutomationML data format. The model is demonstrated on a use-case for disassembling electric vehicle batteries to support their remanufacturing for stationary battery applications.
@@ I Introduction
The significant environmental impact of manufacturing industries underscores the importance of addressing sustainability in the context of recycling and remanufacturing of products [1]. Although the term ”sustainability” is increasingly used in marketing, leading to skepticism and negative perceptions. Distinguishing genuine sustainability efforts from mere marketing strategies is essential. In this paper, we introduce a formalization to model products and production processes, which has the potential to contribute to sustainability specifically at the end of a product life-cycle.
@@
PoPAN offers a structured description of a product, incorporating its components (represented as products), processes, and resources. This approach enables the design of recycling and remanufacturing processes for products that require disassembly for efficient recycling. Furthermore, PoPAN can function as a digital shadow [4], accompanying the product throughout its entire life-cycle and encapsulating all relevant information. It serves as a record of the product’s evolution, including any modifications or alterations it undergoes during its journey, such as missing components like screws. An additional advantage of employing PoPAN for recycling lies in the data collection aspect. Sustainability initiatives begin at the product design phase, where prioritizing recyclability is paramount. By gathering relevant data from recycling processes, manufacturers can analyze and utilize this information to enhance the recyclability of future product designs. This iterative process leads to more sustainable designs and facilitates easier recycling, contributing to overall environmental sustainability efforts.
To facilitate the adoption of this paradigm, the paper also proposes the serialization of the model in the AutomationML [5] data format. This standardized format enhances interoperability and ease of integration within existing industrial frameworks and systems. An important benefit of AutomationML is the advancing integration with the Asset Administration Shell (AAS) [6]. AAS is a set of standards and recommendations to provide interoperability in industrial systems. AAS is typically used on the resource level (e.g., a motor or a robot). On the contrary, the AAS is used for the product level in the presented approach utilizing PoPAN. Moreover, AAS is used for the entire product life-cycle in this approach, including not only the manufacturing phase but also operations, maintenance, and repair, as well as decommissioning/recycling. The paper [7] highlights the use of AAS in the context of product life cycle management.
@@
  • RQ1: How can we adapt the PAN model to primarily reflect the product structure rather than the production process with resource structure?
@@
  • RQ3: To improve the adoption of the proposed approach by industry and academia, how can be the proposed model represented in the AutomationML data format?
The proposed modeling approach is demonstrated in a case study on disassembling electric vehicle (EV) batteries for recycling and remanufacturing [8].
@@ II State of the Art
The proposed approach is built on the top of 3 main domains including (i) Asset Administration Shell, (ii) AutomationML data format, and (iii) Model-Driven Engineering and Product-Process-Resource Asset Network. These research areas are described in detail in the following subsections.
@@ II-A Asset Administration Shell
One of the key standards in the Industry 4.0 initiative is the Asset Administration Shell (AAS) [9, 10]. AAS was introduced by Platform Industry 4.0 as a promising approach to enable interoperability in various industries. It facilitates data sharing between value chain partners, standardizes data security, and establishes technology-neutral semantic standards, making it crucial to achieving Industry 4.0 goals. The AAS enables diverse communication channels and applications, serving as a bridge between physical objects and the digital world [6].
The AAS is a digital representation of an asset and comprises multiple submodels that represent a specific aspect or set of characteristics of an asset. Within a submodel, various submodel elements are defined. Submodel elements are the individual components or entities that make up the submodel. They represent specific properties, operations, parameters, references, files, or other relevant information associated with the asset. The purpose of submodels is to provide a structured approach to organizing and managing the elements within an AAS. By grouping related elements, it becomes easier to define and enforce standards for specific sets of information.
@@ II-B AutomationML
The engineering phase of multi-disciplinary mechatronic systems requires data exchange across software tools belonging to diverse engineering domains (including electric planning, mechanic design, maintenance, etc.). Engineering data can be efficiently captured and exchanged in the data format AutomationML, which is a standardized extension of XML for the interdisciplinary exchange of planning data for production systems and processes. The data format AutomationML is an open, platform-independent standard known as IEC 62714. Detailed information about this standard including various best practices and application recommendations, and how to model specific information, can be found on the Web pages of the AutomationML association [^1] Online: https://www.automationml.org/.
@@
  • CAEX (IEC 62424) – Topology a hierarchy of objects, properties, and relationships among objects
@@
The basis of the object model in AutomationML is CAEX, which is intended for capturing engineering information in object-oriented form. Each object can have attributes and references to other objects (i.e., internal links or mapping objects, and references via object identifiers). It can also include links to other information, represented in COLLADA or PLCopen XML. In CAEX, a model is represented via the following four basic cornerstones. Role Class Libraries are intended to define object semantics. Each object can have more than one role. Interface Class Libraries provide specifications of object interfaces. Classes of objects are modeled as objects of System Unit Class Libraries. Each system unit class poses a generic object prototype, which can be instantiated when describing a specific system. For better interoperability, semantic mapping to terminology defined as ECLASS based on IEC 61360 can be used in a standardized way. The system description itself is represented in AutomationML/CAEX by instance hierarchy. In the tree structure behind the instance hierarchy root element, Internal Elements, and their sub-elements are defined. Internal elements in most cases reference their prototypes in the frame of the System Unit Class, supported roles from Role Class Libraries, and interfaces from the interface class libraries. Topological relationships are represented as Internal Links.
AutomationML is one of the most complex data formats for system engineering, which is suitable for a wide range of tasks, including the representation of the Asset Administration Shell [11, 12]. Such an AutomationML representation can be imported to the AASX format [9] in the tool AASX Package Explorer [^2] Online: https://github.com/eclipse-aaspe. The application of AutomationML in the end-of-life phase is addressed in [13, 14], which are focused on production systems and resources, whereas this paper is mainly focused on products together with production processes.
@@ II-C Model-Driven Engineering and Product-Process-Resource Asset Network
Model-driven engineering (MDE) highlights the role of models as core artifacts for engineering and integration development. Similarly, like object-orientation in object-oriented programming, where everything is represented by objects (with specific attributes and methods), in model-driven engineering, everything is a model [15]. Initial efforts in MDE started with the use of the Meta Object Facility and the Unified Modeling Language (UML) [16]. MDE is highly efficient for translating information between various representations (i.e., domain models or data formats) [17, 18]; and it provides suitable support for multi-disciplinary engineering projects [19]. One of the plausible models to support the model-driven engineering is a Product-Process-Resource model [20], categorizing artifacts relevant for industrial production into these three categories including mappings or links among these categories.
The Product-Process-Resource Asset Network (PAN) [3] is a graph-based formalism for expressing relationships among products, production processes, and resources. In the domain of Cyber-Physical Production Systems (CPPSs), dependencies among products, processes, and resources are frequently left implicit. The PAN addresses this gap by providing a structured framework for explicitly articulating and overseeing the dependencies among product, process, and resource assets. The PAN modeling is suitable for capturing and sharing knowledge among engineers and respective engineering tools and for identifying changes during engineering processes of multi-disciplinary system engineering projects. PAN usage optimizes engineering processes’ efficiency and enhances operations by modeling manufacturing processes, tracking resource usage, and monitoring asset performance. Through data integration, visualization, and analysis, PAN enables informed decision-making [21].
@@ III Product-oriented Product–Process–Resource Asset Network (PoPAN)
Figure 1: The proposed PoPAN model for a generic product (on the right), which has been created based on the Product Structure model (on the left). The PoPAN graph is substantiated by products, which have assigned processes, and the edges (their types, orientation, and arrow types) have a crucial role.
@@
Recycling and remanufacturing products pose challenges due to limited engineering data availability. While manufacturing processes are well-specified and documented, the absence of detailed data for a product in its end-of-life phase complicates recycling and remanufacturing efforts. Leveraging the PAN approach, which has proven efficiency in product manufacturing, presents a promising solution. However, traditional PAN methods describe one specific assembly process, which may not capture the entire variability of assembly and disassembly operations. To address this limitation, this paper proposes extending the PAN paradigm to describe products structurally. By doing so, multiple disassembly processes can be generated from a single description, accommodating the diverse disassembly options frequently encountered in practice. This structural representation enhances flexibility and efficiency in creating recycling and remanufacturing processes.
@@
To create a product description that offers a comprehensive understanding of the product’s structure, including the specific processes necessary for both assembly and disassembly, along with the resources required for each process, the initial step is to construct a Product Structure graph depicting dependencies between product’s components (see Fig. 1 on the left). Since the PPR approach is used, to preserve the name of elements, the product’s components are depicted also as products. Within this framework, three types of products are distinguished (1) Elementary Product: a product that cannot be further disassembled, (2) SubProduct: a product that is considered elementary in terms of assembly/disassembly procedures, but is composed of several parts and can be further disassembled, and (3) Fastener Product: a product that serves as a fastener for other products (e.g., screws). The graph delineates two types of edges connecting individual products, forming the overall structure. The solid line edge illustrates physical continuity between products and the dashed line edge, on the other hand, forms a structural continuity between the products, i.e., dependencies that indicate certain products cannot be assembled or disassembled until others are. Moreover, both types of edges end with arrows on both ends representing the assembly/disassembly direction. Open arrows represent assembly direction, full arrows represent disassembly direction, and Initial Product and Last Product are highlighted for better orientation in the graph. For example, when disassembling Product 3, it is essential to consider all three structural dependencies. Considering the orientation by the full arrows, Product 1 and Product 2 are not blocking Product 3 from being disassembled. However, the Last Product is blocking the disassembly of Product 3, meaning that the Last Product must be disassembled before Product 3. Additionally, frequently used terms final product and semi-product in the sense of a classic PAN are highlighted to show intermediate steps in an assembly operation. However, the PoPAN approach diverges from relying on these terms.
@@
Expanding the Product Structure graph with the PAN/PPR approach involves incorporating processes and resources alongside products. This structured visualization facilitates understanding the relationships between products and corresponding processes, which are supported by the required resources. As depicted in Fig. 1 on the right, in the extended Product Structure graph called Product-oriented Product–Process–Resource Asset Network (PoPAN), each product is connected with a corresponding process by a simple line edge or by a simple line edge with a question mark above it. The question mark indicates conditional linking between the product and process, where the process sequence might not correspond with the structure of a product, which is then resolved by a red dotted line edge determining the sequence of the processes. Lastly, the graph contains green circle marks with a letter which depict edges connecting processes with their required resources. As with the previous Product Structure graph, the arrows form the sense of orientation in the graph and the connections between products, processes, and resources form the edges.
@@
    Upon arriving at Fastener Product i, the conditional linking product to the process edge is examined.
@@
    Returning to Fastener Product i, the physical continuity of the product’s edge is verified in the correct direction, leading to Product 1.
@@
    Upon reaching Product 1, the structural continuity of the product’s edge is examined, indicating the Product 3 is to be assembled subsequent to Product 1.
@@
    Fastener Process i is noted with Fastener Product i simultaneously with Resource 1.
As outlined in the preceding steps, when assembling a product the initial edge to examine, if present, is the structural continuity of the product’s edge. This edge determines whether the product can be assembled before other products. Once the feasibility of assembly/disassembly is established, the next edges to consider are the linking products to the process edge and conditional linking products to the process edge, which assign processes to products. Additionally, for the conditional linking products to the process edge, the sequence of the process edge is presented which ensures the processes are sequenced correctly. Afterward, the linking product to the process edge is followed to assign a resource to a product after which the physical continuity of the product’s edge can be taken to continue to the next product where the whole process is repeated.
These rules, which specify the order in which edges should be traversed along with the assembly/disassembly direction, enable PoPAN to determine the correct path for creating assembly/disassembly operations. This capability facilitates the efficient design of recycling or remanufacturing processes of products.